Amanda Rice has just arranged to purchase a $560,000 vacation home in the Bahamas with a 25 percent down payment. The mortgage has a 6.2 percent APR compounded monthly and calls for equal monthly payments over the next 30 years. Her first payment will be due one month from now. However, the mortgage has an eight-year balloon payment, meaning that the balance of the loan must be paid off at the end of Year 8. There were no other transaction costs or finance charges.

How much will Amanda’s balloon payment be in eight years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Amanda's Balloon Payment should be equal to the portion of the original loan amount, outstanding at the end of Year 8. The same can be determined by calculating the total present value (at end of Year 8) of equal monthly payments remaining after year 8

House Price = $ 560000, Down Payment = 25 %, Loan Amount = (1-0.25) x 560000 = $ 420000

Interest Rate = 6.2 % APR or (6.2 / 12) ~ 0.5167 % per month

Tenure = 30 years or 360 months

Let the equal monthly payments be $ K

Therefore, 420000 = K x (1/0.005167) x [1-{1/(1.005167)^(360)}]

K = $ 2572.48

Balloon Payment = 2572.48 x (1/0.005167) x [1-{1/(1.005167)^(264)}] = $ 370157.82